Please help, I'm not sure how to solve this word problem:
"Mr.X" bowled an average score of 150 points per game for 12 games. What average score must he bowl for the next 6 games to raise his average score per game for the next 18 games to 160?

Answer by richwmiller(17219)

(12*150+6y)/18=160
1800/18+6y/18=160
100+y/3=160
y/3=60
y=180
check

(12*150+180*6)/18=160
(1800+1080)/18=160
2880/18=160
160=160
ok